OpenStudy (kkutie7):

Any one good at titration? If 0.4905 g of pure KHP solid were titrated with NaOH and the endpoint was reached at 28.08 mL, what is the concentration of the NaOH solution? (Note: MW of KHP = 204.223 g/mol).

OpenStudy (aaronq):

You need the balanced equation, do you have that?

OpenStudy (kkutie7):

Feel like looking over my work when you get the chance? =) \[.4905g*\frac{ 1mol }{204.223g }=.0024mol\] \[28.08ml*\frac{ 1L }{ 1000ml }=.028L\] \[\frac{ .0024 }{ .028 }=.00857M\]

OpenStudy (kkutie7):

No equation was provided.

OpenStudy (kkutie7):

@aaronq

OpenStudy (aaronq):

yeah it's right, good job!

OpenStudy (kkutie7):

thank you